# On multi-graded Proj schemes

### v4

@@ 4 Examples arising from reductive algebraic groups
In this section we show that a number of varieties of interest in Geometric Representation Theory can be described as Proj schemes associated with certain graded rings in a canonical way (which, in particular, does not require a specific choice of dominant character). We see these examples as a motivation to study the general theory considered in the rest of the paper, but none of the considerations of the present sections will be used in proofs of results of other sections. (In Section 5, these results will only be used as illustrations of certain general statements.)
@@ 4.1 Flag varieties
We will give two independent proofs of Proposition 4.2. The first one uses the description of the Proj-scheme as a geometric quotient under the action of a diagonalizable group scheme recalled in §3.2. The second one uses results of [BS03] on ample families of line bundles.
@@
and composition with the surjection $\widetilde{\mathbf{T}}\to\mathbf{T}$ induces an embedding $\mathbb{X}\hookrightarrow\widetilde{\mathbb{X}}$ whose cokernel is finite. For $\lambda\in\mathbb{X}$, under the identification (4.5) the line bundle on $\mathbf{G}/\mathbf{B}$ associated with $\lambda$ identifies with the line bundle on $\widetilde{\mathbf{G}}/\widetilde{\mathbf{B}}$ associated with its image in $\widetilde{\mathbb{X}}$; hence $A$ identifies with the subring of $\widetilde{A}$ given by the sum of the components whose label is in $\mathbb{X}$. In view of Proposition 3.20, we deduce an isomorphism
@@
$\left(\bigotimes_{\alpha\in\mathfrak{R}_{\mathrm{s}}}\mathrm{S}(A_{\varpi_{\alpha}})\right)\to A$ (4.7)
@@
Finally, choose a numbering $\alpha_{1},\dots,\alpha_{r}$ of the elements in $\mathfrak{R}_{\mathrm{s}}$. It is a standard fact (see e.g. [Ja03, Proposition II.4.4]) that ${\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\lambda)$ is ample for any $\lambda\in\mathbb{X}_{++}$; in particular, ${\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\varpi_{\alpha_{1}}+\cdots+\varpi_{\alpha_{r}})$ is ample. Comparing the characterization of ample line bundles in [StP, Tag 01PS] with [BS03, Proposition 1.1] we deduce that the collection
@@
$\bigsqcup_{\alpha\in\mathfrak{R}_{\mathrm{s}}}I_{\alpha}\to\mathfrak{R}_{\mathrm{s}};$
@@ 4.2 Some vector bundles over the flag variety
i.e. the quotient of the product $\mathbf{G}\times V$ by the (free) action of $\mathbf{B}$ defined by $b\cdot(g,x)=(gb^{-1},b\cdot x)$. (This construction is a special case of that considered in [Ja03, §I.5.14].) It is a vector bundle over $\mathbf{G}/\mathbf{B}$ (in particular, a smooth variety). For $\lambda\in\mathbb{X}$ we will denote by ${\mathcal{O}}_{\mathbf{G}\times^{\mathbf{B}}V}(\lambda)$ the pullback to $\mathbf{G}\times^{\mathbf{B}}V$ of the line bundle ${\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\lambda)$. By (4.1), for $\lambda,\mu\in\mathbb{X}$ we have a canonical isomorphism
@@
which provides a structure of $\mathbb{X}$-graded ring on
@@
${\mathcal{O}}(\widetilde{V})\otimes A_{\lambda}\to(A_{V})_{\lambda}.$ (4.11)
@@
$\mathbf{G}\times^{\mathbf{B}}V\to\widetilde{V}.$ (4.12)
@@
Now $\mathcal{V}^{\bot}$ is the vector bundle on $\mathbf{G}/\mathbf{B}$ associated with the orthogonal $V^{\bot}\subset\widetilde{V}^{*}$ of $V\subset\widetilde{V}$. Hence, if we denote by $\Lambda\subset\mathbb{X}$ the (finite) subset consisting of the $\mathbf{T}$-weights appearing in the various exterior powers of $V^{\bot}$, then each $\wedge^{i}\mathcal{V}^{\bot}$ is an iterated extension of line bundles of the form ${\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\mu)$ with $\mu\in\Lambda$. If we choose $N$ such that $\langle\mu,\alpha^{\vee}\rangle\geqslant-N$ for any $\mu\in\Lambda$ and any simple root $\alpha$, then if $\lambda\in\mathbb{X}_{+}$ satisfies $\langle\lambda,\alpha^{\vee}\rangle\geqslant N$ for any simple root $\alpha$, each weight $\lambda+\mu$ with $\mu\in\Lambda$ is dominant. Since line bundles on $\mathbf{G}/\mathbf{B}$ associated with dominant weights have no higher cohomology (this is Kempf’s vanishing theorem, see [Ja03, Proposition II.4.5]), for such $\lambda$ we have
@@
${\mathcal{O}}(\widetilde{V})\otimes A\to A_{V}.$ (4.15)
@@
$\mathbf{G}\times^{\mathbf{B}}V\to(\mathbf{G}\times^{\mathbf{B}}V)_{\mathrm{aff}}$
@@
Finally, we observe that if $\lambda\in\mathbb{X}_{++}$ the line bundle ${\mathcal{O}}_{\mathbf{G}\times^{\mathbf{B}}V}(\lambda)$ is ample by [StP, Tag 0892] and the similar claim for $\mathbf{G}/\mathbf{B}$. As in the second proof of Proposition 4.2, these arguments show that the conditions in [BS03, Corollary 4.6] are satisfied, and this statement provides the desired identification of $\mathbf{G}\times^{\mathbf{B}}V$ with ${\operatorname{Proj}}^{\mathbb{X}}(A_{V})$. ∎
@@ 4.3 The Springer resolution
${\mathcal{O}}(\mathbf{g})\otimes A\to A_{\mathbf{u}}^{\prime}\to A_{\mathbf{u}},$
@@ 5.1 Sheaves associated with graded modules
$\mathsf{L}:{\mathsf{Mod}}^{M}(A)/{\mathsf{Mod}}^{M}(A)_{\mathrm{neg}}\to\mathsf{QCoh}({\operatorname{Proj}}^{M}(A)).$
@@
$Q_{0}\to\Gamma\bigl{(}{\operatorname{Proj}}^{M}(A),\widetilde{Q}\bigr{)}$
@@
commutes, where the upper horizontal arow is the natural map, and the vertical ones are induced by the morphisms from Fact 5.2.
@@ 5.2 Twisting sheaves
We now drop the assumption that $A$ is noetherian. Note that $\mathcal{O}_{\mathrm{Proj}^{M}(A)}(\alpha)$ is not a line bundle in general, even in the $\mathbb{N}$-graded setting. If $\alpha,\alpha^{\prime}\in M$, since $A(\alpha)\otimes_{A}A(\alpha^{\prime})=A(\alpha+\alpha^{\prime})$, Proposition 5.3 implies that there is a canonical map
@@
a structure of $M$-graded sheaf of $\mathcal{O}_{\mathrm{Proj}^{M}(A)}$-algebras, and an $M$-graded ring structure on
@@
a structure of $M$-graded sheaf of modules over (5.2), and on
@@
a structure of $M$-graded module over (5.3).
@@ 5.3 Graded modules associated with sheaves
a structure of $M$-graded module over the ring (5.3), hence over the ring $A$ by Fact 5.7(1). Denoting by ${\mathsf{Mod}}({\mathcal{O}}_{\mathrm{Proj}^{M}(A)})$ the abelian category of sheaves of ${\mathcal{O}}_{\mathrm{Proj}^{M}(A)}$-modules, this construction provides a functor
$\Gamma_{\bullet}:{\mathsf{Mod}}({\mathcal{O}}_{\mathrm{Proj}^{M}(A)})\to{\mathsf{Mod}}^{M}(A).$
@@
    $\widetilde{Q}(\alpha)_{|D_{\dagger}(S)}\to\widetilde{Q(\alpha)}_{|D_{\dagger}(S)}$
@@
    ${\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\alpha)_{|D_{\dagger}(S)}\otimes_{{\mathcal{O}}_{D_{\dagger}(S)}}{\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\alpha^{\prime})_{|D_{\dagger}(S)}\to{\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\alpha+\alpha^{\prime})_{|D_{\dagger}(S)}$
@@
    ${\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\alpha)_{|D_{\dagger}(S)}\otimes_{{\mathcal{O}}_{D_{\dagger}(S)}}{\mathcal{Q}}(\alpha^{\prime})|_{D_{\dagger}(S)}\to{\mathcal{Q}}(\alpha+\alpha^{\prime})_{|D_{\dagger}(S)}$
@@
Since $\alpha\in M[\underline{S}]$, there exists an invertible element $a\in A_{S}$ of degree $\alpha$. Then the map $x\mapsto a\cdot x$ induces an isomorphism of $A_{(S)}$-modules $A_{(S)}\cong(A(\alpha))_{(S)}=(A_{S})_{\alpha}$. We deduce (1). Similarly, given a graded $A$-module $Q$, the map $q\mapsto a\cdot q$ induces an isomorphism of $A_{(S)}$-modules $Q_{(S)}\cong(Q(\alpha))_{(S)}=(Q_{S})_{\alpha}$, which implies (2). The statement in (3) is the special case of (2) where $Q=A(\alpha^{\prime})$. Finally, (4) is a direct consequence of (3). ∎
@@
Now an element of $\Gamma_{\bullet}({\mathcal{Q}})_{(S)}$ can be represented by a fraction $\frac{m}{s}$ where $s\in S$ and $\in\Gamma_{\bullet}({\mathcal{Q}})_{\deg(s)}$, i.e. $m$ is a global section of ${\mathcal{Q}}(\deg(s))$. Consider
@@
wjere the isomorphism is provided by Proposition 5.8(4).
@@ 5.4 Maximally relevant families
    Assume that $M=\mathbb{Z}$, and that $A$ is generated by $A_{1}$ as an $A_{0}$-algebra. (In particular, we are in the $\mathbb{N}$-graded case.) If $(a_{i}:i\in I)$ is a family of generators of $A_{1}$ as an $A_{0}$-module, since any non nilpotent relevant element must have positive degree the condition of Corollary 3.10 is satisfied. Clearly, each $a_{i}$ is maximally relevant.
@@
Assume that (5.8) is satisfied. Then for any $\alpha\in M$, the quasi-coherent ${\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}$-module ${\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\alpha)$ is a line bundle. Moreover:
@@
In the setting of §4.1, for $\lambda\in\mathbb{X}$ the line bundle ${\mathcal{O}}_{{\operatorname{Proj}}^{\mathbb{X}}(A)}(\lambda)$ corresponds to the line bundle ${\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\lambda)$. A similar comment applies in the setting considered in §4.2.
@@
${\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})\to{\operatorname{Proj}}^{M}(A)$
@@
Since ${\operatorname{Proj}}^{M}(A)$ is covered by the affine open subschemes $(D_{\dagger}(S):S\in{\mathcal{F}})$, we have $\widetilde{Q}=0$ iff $Q_{(S)}=0$ for any $S\in{\mathcal{F}}$. Now since ${\mathcal{F}}\subset{\mathcal{F}}^{\mathrm{max}}$, for any $S\in{\mathcal{F}}$ the ring $A_{S}$ has invertible elements of all degrees, so that $Q_{(S)}=0$ iff $Q_{S}=0$. Finally, it is clear from definitions that $Q_{S}=0$ iff for any $q\in Q$ there exists $s\in S$ such that $s\cdot q=0$. ∎
@@
    First, assume that $A_{0}$ is noetherian, that $A$ is generated by $A_{1}$ as an $A_{0}$-algebra, and moreover that $A_{1}$ is finite as an $A_{0}$-module. (This is the setting considered in [StP, Tag 01YR].) For $N\in\mathbb{Z}$ we set $A_{\geqslant N}=\bigoplus_{m\geqslant N}A_{m}$. Then $Q$ is negligible iff for any $q\in Q$ there exists $N>0$ such that $A_{\geqslant N}\cdot q=0$. (Since $A_{\geqslant N}=(A_{\geqslant 1})^{N}$, this condition is equivalent to $Q$ being $A_{\geqslant 1}$-power torsion in the sense of [StP, Tag 05E6].) In fact, choose $(a_{i}:i\in I)$ as in Remark 5.10, with $I$ finite; then we can choose =${\mathcal{F}}=\{(a_{i})^{\mathbb{N}}:i\in I\}$. If $Q$ satisfies our condition, then it is negligible by Lemma 5.15. On the other hand, assume that $Q$ is negligible, and let $q\in Q$. By the lemma, for any $i\in I$ there exists $n_{i}\in\mathbb{Z}_{>0}$ such that $(a_{i})^{n_{i}}\cdot q=0$. Then if $N=1+\sum_{i\in I}(n_{i}-1)$, we have
@@
    Now, consider the setting of §4.1, and recall the notation of Remarks 4.7 and 5.10. For $N\in\mathbb{Z}$ we denote by $\mathbb{X}_{\geqslant N}\subset\mathbb{X}$ the submonoid consisting of elements $\lambda$ which satisfy $\langle\lambda,\alpha^{\vee}\rangle\geqslant N$ for any simple root $\alpha$. (In particular, we have $\mathbb{X}_{\geqslant 0}=\mathbb{X}_{+}$, and $\mathbb{X}_{\geqslant 1}=\mathbb{X}_{++}$.) We also denote by $A_{\geqslant N}$ the sum of the components in $A$ whose degrees belong to $\mathbb{X}_{\geqslant N}$. We will use similar notation for $\widetilde{\mathbb{X}}$ and $\widetilde{A}$. Then $Q$ is negligible iff for any $q\in Q$ there exists $N>0$ such that $A_{\geqslant N}\cdot q=0$. In fact, in Lemma 5.15 we can take =${\mathcal{F}}=\{(g_{\sigma})^{\mathbb{N}}:\sigma\in E\}$. Since each $g_{\sigma}$ belongs to $A_{\geqslant 1}$, it is clear from this lemma that if $Q$ satisfies our condition, then it is negligible. On the other hand assume that $Q$ is negligible, and fix $q\in Q$. By the lemma, for any $\sigma\in E$ there exists $N_{\sigma}$ such that $(g_{\sigma})^{N_{\sigma}}\cdot q=0$. If $N=1+\sum_{\sigma}(N_{\sigma}n_{\sigma}-1)$, we have
@@
    Consider the setting of §4.2. Defining $\mathbb{X}_{\geqslant N}$ as above, and then $(A_{V})_{\geqslant N}$ as the sum of the graded components of $A_{V}$ whose label belongs to $\mathbb{X}_{\geqslant N}$, one checks using Lemma 4.11 and (5.9) that a graded $A_{V}$-module $Q$ is negligible iff for any $q\in Q$ there exists $N>0$ such that $(A_{V})_{\geqslant N}\cdot q=0$. Similar comments apply in the setting of §4.3.
@@
${\mathsf{Mod}}^{M}(A)\xrightarrow{Q\mapsto\widetilde{Q}}\mathsf{QCoh}({\operatorname{Proj}}^{M}(A))\to{\mathsf{Mod}}({\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)})$
@@
To prove the proposition we need to define functorial morphisms $\varepsilon_{{\mathcal{Q}}}:\widetilde{\Gamma_{\bullet}({\mathcal{Q}})}\to{\mathcal{Q}}$ for ${\mathcal{Q}}\in{\mathsf{Mod}}({\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)})$ and $\eta_{Q}:Q\to\Gamma_{\bullet}(\widetilde{Q})$ for $Q\in{\mathsf{Mod}}^{M}(A)$, which satisfy the usual zigzag relations. Here $\varepsilon$ is provided by Lemma 5.9. (This does not require any assumption.) To define $\eta$, we observe that for $Q\in{\mathsf{Mod}}^{M}(A)$ we have

### v5

@@ 4 Examples arising from reductive algebraic groups
In this section we show that a number of varieties of interest in Geometric Representation Theory can be described as Proj schemes associated with certain graded rings in a canonical way (which, in particular, does not require a specific choice of a dominant character). We see these examples as a motivation to study the general theory considered in the rest of the paper, but none of the considerations of the present sections will be used in proofs of results of other sections. (In Section 5, these results will only be used as illustrations of certain general statements.)
@@ 4.1 Flag varieties
We will give two independent proofs of Proposition 4.2. The first one uses the description of the Proj-scheme as a geometric quotient under the action of a diagonalizable group scheme recalled in §3.2. The second one uses results of [BS03] on ample families of invertible sheaves.
@@
and composition with the surjection $\widetilde{\mathbf{T}}\to\mathbf{T}$ induces an embedding $\mathbb{X}\hookrightarrow\widetilde{\mathbb{X}}$ whose cokernel is finite. For $\lambda\in\mathbb{X}$, under the identification (4.5) the invertible sheaf on $\mathbf{G}/\mathbf{B}$ associated with $\lambda$ identifies with the invertible sheaf on $\widetilde{\mathbf{G}}/\widetilde{\mathbf{B}}$ associated with its image in $\widetilde{\mathbb{X}}$; hence $A$ identifies with the subring of $\widetilde{A}$ given by the sum of the components whose label is in $\mathbb{X}$. In view of Proposition 3.22, we deduce an isomorphism
@@
$\left(\bigotimes_{\alpha\in\mathfrak{R}_{\mathrm{s}}}\mathrm{S}(A_{\varpi_{\alpha}})\right)\longrightarrow A$ (4.7)
@@
Finally, choose a numbering $\alpha_{1},\dots,\alpha_{r}$ of the elements in $\mathfrak{R}_{\mathrm{s}}$. It is a standard fact (see e.g. [Ja03, Proposition II.4.4]) that ${\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\lambda)$ is ample for any $\lambda\in\mathbb{X}_{++}$; in particular, ${\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\varpi_{\alpha_{1}}+\cdots+\varpi_{\alpha_{r}})$ is ample. Comparing the characterization of ample invertible sheaves in [StP, Tag 01PS] with [BS03, Proposition 1.1] we deduce that the collection
@@
$\bigsqcup_{\alpha\in\mathfrak{R}_{\mathrm{s}}}I_{\alpha}\longrightarrow\mathfrak{R}_{\mathrm{s}};$
@@ 4.2 Some vector bundles over the flag variety
i.e. the quotient of the product $\mathbf{G}\times V$ by the (free) action of $\mathbf{B}$ defined by $b\cdot(g,x)=(gb^{-1},b\cdot x)$. (This construction is a special case of that considered in [Ja03, §I.5.14].) It is a vector bundle over $\mathbf{G}/\mathbf{B}$ (in particular, a smooth variety). For $\lambda\in\mathbb{X}$ we will denote by ${\mathcal{O}}_{\mathbf{G}\times^{\mathbf{B}}V}(\lambda)$ the pullback to $\mathbf{G}\times^{\mathbf{B}}V$ of the invertible sheaf ${\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\lambda)$. By (4.1), for $\lambda,\mu\in\mathbb{X}$ we have a canonical isomorphism
@@
which provides a structure of a $\mathbb{X}$-graded ring on
@@
${\mathcal{O}}(\widetilde{V})\otimes A_{\lambda}\longrightarrow(A_{V})_{\lambda}.$ (4.11)
@@
$\mathbf{G}\times^{\mathbf{B}}V\longrightarrow\widetilde{V}.$ (4.12)
@@
Now $\mathcal{V}^{\bot}$ is the vector bundle on $\mathbf{G}/\mathbf{B}$ associated with the orthogonal $V^{\bot}\subset\widetilde{V}^{*}$ of $V\subset\widetilde{V}$. Hence, if we denote by $\Lambda\subset\mathbb{X}$ the (finite) subset consisting of the $\mathbf{T}$-weights appearing in the various exterior powers of $V^{\bot}$, then each $\wedge^{i}\mathcal{V}^{\bot}$ is an iterated extension of invertible sheaves of the form ${\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\mu)$ with $\mu\in\Lambda$. If we choose $N$ such that $\langle\mu,\alpha^{\vee}\rangle\geqslant-N$ for any $\mu\in\Lambda$ and any simple root $\alpha$, then if $\lambda\in\mathbb{X}_{+}$ satisfies $\langle\lambda,\alpha^{\vee}\rangle\geqslant N$ for any simple root $\alpha$, each weight $\lambda+\mu$ with $\mu\in\Lambda$ is dominant. Since invertible sheaves on $\mathbf{G}/\mathbf{B}$ associated with dominant weights have no higher cohomology (this is Kempf’s vanishing theorem, see [Ja03, Proposition II.4.5]), for such $\lambda$ we have
@@
${\mathcal{O}}(\widetilde{V})\otimes A\longrightarrow A_{V}.$ (4.15)
@@
$\mathbf{G}\times^{\mathbf{B}}V\longrightarrow(\mathbf{G}\times^{\mathbf{B}}V)_{\mathrm{aff}}$
@@
Finally, we observe that if $\lambda\in\mathbb{X}_{++}$ the invertible sheaf ${\mathcal{O}}_{\mathbf{G}\times^{\mathbf{B}}V}(\lambda)$ is ample by [StP, Tag 0892] and the similar claim for $\mathbf{G}/\mathbf{B}$. As in the second proof of Proposition 4.2, these arguments show that the conditions in [BS03, Corollary 4.6] are satisfied, and this statement provides the desired identification of $\mathbf{G}\times^{\mathbf{B}}V$ with ${\operatorname{Proj}}^{\mathbb{X}}(A_{V})$. ∎
@@ 4.3 The Springer resolution
${\mathcal{O}}(\mathbf{g})\otimes A\longrightarrow A_{\mathbf{u}}^{\prime}\to A_{\mathbf{u}},$
@@ 5.1 Sheaves associated with graded modules
$\mathsf{L}:{\mathsf{Mod}}^{M}(A)/{\mathsf{Mod}}^{M}(A)_{\mathrm{neg}}\longrightarrow\mathsf{QCoh}({\operatorname{Proj}}^{M}(A)).$
@@
$Q_{0}\longrightarrow\Gamma\bigl{(}{\operatorname{Proj}}^{M}(A),\widetilde{Q}\bigr{)}$
@@
commutes, where the upper horizontal arrow is the natural map, and the vertical ones are induced by the morphisms from Fact 5.2.
@@ 5.2 Twisting sheaves
We now drop the assumption that $A$ is noetherian. Note that $\mathcal{O}_{\mathrm{Proj}^{M}(A)}(\alpha)$ is not an invertible sheaf in general, even in the $\mathbb{N}$-graded setting. If $\alpha,\alpha^{\prime}\in M$, since $A(\alpha)\otimes_{A}A(\alpha^{\prime})=A(\alpha+\alpha^{\prime})$, Proposition 5.3 implies that there is a canonical map
@@
a structure of an $M$-graded sheaf of $\mathcal{O}_{\mathrm{Proj}^{M}(A)}$-algebras, and an $M$-graded ring structure on
@@
a structure of an $M$-graded sheaf of modules over (5.2), and on
@@
a structure of an $M$-graded module over (5.3).
@@ 5.3 Graded modules associated with sheaves
a structure of an $M$-graded module over the ring (5.3), hence over the ring $A$ by Fact 5.7(1). Denoting by ${\mathsf{Mod}}({\mathcal{O}}_{\mathrm{Proj}^{M}(A)})$ the abelian category of sheaves of ${\mathcal{O}}_{\mathrm{Proj}^{M}(A)}$-modules, this construction provides a functor
$\Gamma_{\bullet}:{\mathsf{Mod}}({\mathcal{O}}_{\mathrm{Proj}^{M}(A)})\longrightarrow{\mathsf{Mod}}^{M}(A).$
@@
    $\widetilde{Q}(\alpha)_{|D_{\dagger}(S)}\longrightarrow\widetilde{Q(\alpha)}_{|D_{\dagger}(S)}$
@@
    $\left({\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\alpha)_{|D_{\dagger}(S)}\right)\otimes_{{\mathcal{O}}_{D_{\dagger}(S)}}\left({\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\alpha^{\prime})_{|D_{\dagger}(S)}\right)\longrightarrow{\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\alpha+\alpha^{\prime})_{|D_{\dagger}(S)}$
@@
    $\left({\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\alpha)_{|D_{\dagger}(S)}\right)\otimes_{{\mathcal{O}}_{D_{\dagger}(S)}}\left({\mathcal{Q}}(\alpha^{\prime})|_{D_{\dagger}(S)}\right)\longrightarrow{\mathcal{Q}}(\alpha+\alpha^{\prime})_{|D_{\dagger}(S)}$
@@
Since $\alpha\in M[\underline{S}]$, there exists an invertible element $a\in A_{S}$ of degree $\alpha$ (see Proposition 2.5(1)). Then the map $x\mapsto a\cdot x$ induces an isomorphism of $A_{(S)}$-modules $A_{(S)}\cong(A(\alpha))_{(S)}=(A_{S})_{\alpha}$. We deduce (1). Similarly, given a graded $A$-module $Q$, the map $q\mapsto a\cdot q$ induces an isomorphism of $A_{(S)}$-modules $Q_{(S)}\cong(Q(\alpha))_{(S)}=(Q_{S})_{\alpha}$, which implies (2). The statement in (3) is the special case of (2) where $Q=A(\alpha^{\prime})$. Finally, (4) is a direct consequence of (3). ∎
@@
Now an element of $\Gamma_{\bullet}({\mathcal{Q}})_{(S)}$ can be represented by a fraction $\frac{m}{s}$ where $s\in S$ and $m\in\Gamma_{\bullet}({\mathcal{Q}})_{\deg(s)}$, i.e. $m$ is a global section of ${\mathcal{Q}}(\deg(s))$. Consider
@@
where the isomorphism is provided by Proposition 5.8(4).
@@ 5.4 Maximally relevant families
    Assume that $M=\mathbb{Z}$, and that $A$ is generated by $A_{1}$ as an $A_{0}$-algebra. If $(a_{i}:i\in I)$ is a family of generators of $A_{1}$ as an $A_{0}$-module, since any non nilpotent relevant element must have positive degree the condition of Corollary 3.11 is satisfied. Clearly, each $a_{i}$ is maximally relevant.
@@
Assume that (5.8) is satisfied. Then for any $\alpha\in M$, the quasi-coherent ${\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}$-module ${\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\alpha)$ is an invertible sheaf. Moreover:
@@
In the setting of §4.1, for $\lambda\in\mathbb{X}$ the invertible sheaf ${\mathcal{O}}_{{\operatorname{Proj}}^{\mathbb{X}}(A)}(\lambda)$ corresponds to the invertible sheaf ${\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\lambda)$. A similar comment applies in the setting considered in §4.2.
@@
${\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})\longrightarrow{\operatorname{Proj}}^{M}(A)$
@@
Since ${\operatorname{Proj}}^{M}(A)$ is covered by the affine open subschemes $(D_{\dagger}(S):S\in{\mathcal{F}})$, we have $\widetilde{Q}=0$ if and only if $Q_{(S)}=0$ for any $S\in{\mathcal{F}}$. Now since ${\mathcal{F}}\subset{\mathcal{F}}^{\mathrm{max}}$, for any $S\in{\mathcal{F}}$ the ring $A_{S}$ has invertible elements of all degrees, so that $Q_{(S)}=0$ if and only if $Q_{S}=0$. Finally, it is clear from the definitions that $Q_{S}=0$ if and only if for any $q\in Q$ there exists $s\in S$ such that $s\cdot q=0$. ∎
@@
    First, assume that $A_{0}$ is noetherian, that $A$ is generated by $A_{1}$ as an $A_{0}$-algebra, and moreover that $A_{1}$ is finite as an $A_{0}$-module. (This is the setting considered in [StP, Tag 01YR].) For $N\in\mathbb{Z}$ we set $A_{\geqslant N}=\bigoplus_{m\geqslant N}A_{m}$. Then $Q$ is negligible if and only if for any $q\in Q$ there exists $N>0$ such that $A_{\geqslant N}\cdot q=0$. (Since $A_{\geqslant N}=(A_{\geqslant 1})^{N}$, this condition is equivalent to $Q$ being $A_{\geqslant 1}$-power torsion in the sense of [StP, Tag 05E6].) In fact, choose $(a_{i}:i\in I)$ as in Remark 5.10, with $I$ finite; then we can choose =${\mathcal{F}}=\{(a_{i})^{\mathbb{N}}:i\in I\}$. If $Q$ satisfies our condition, then it is negligible by Lemma 5.15. On the other hand, assume that $Q$ is negligible, and let $q\in Q$. By the lemma, for any $i\in I$ there exists $n_{i}\in\mathbb{Z}_{>0}$ such that $(a_{i})^{n_{i}}\cdot q=0$. Then if $N=1+\sum_{i\in I}(n_{i}-1)$, we have
@@
    Now, consider the setting of §4.1, and recall the notation of Remarks 4.7 and 5.10. For $N\in\mathbb{Z}$ we denote by $\mathbb{X}_{\geqslant N}\subset\mathbb{X}$ the submonoid consisting of elements $\lambda$ which satisfy $\langle\lambda,\alpha^{\vee}\rangle\geqslant N$ for any simple root $\alpha$. (In particular, we have $\mathbb{X}_{\geqslant 0}=\mathbb{X}_{+}$, and $\mathbb{X}_{\geqslant 1}=\mathbb{X}_{++}$.) We also denote by $A_{\geqslant N}$ the sum of the components in $A$ whose degrees belong to $\mathbb{X}_{\geqslant N}$. We will use similar notation for $\widetilde{\mathbb{X}}$ and $\widetilde{A}$. Then $Q$ is negligible if and only if for any $q\in Q$ there exists $N>0$ such that $A_{\geqslant N}\cdot q=0$. In fact, in Lemma 5.15 we can take =${\mathcal{F}}=\{(g_{\sigma})^{\mathbb{N}}:\sigma\in E\}$. Since each $g_{\sigma}$ belongs to $A_{\geqslant 1}$, it is clear from this lemma that if $Q$ satisfies our condition, then it is negligible. On the other hand assume that $Q$ is negligible, and fix $q\in Q$. By the lemma, for any $\sigma\in E$ there exists $N_{\sigma}$ such that $(g_{\sigma})^{N_{\sigma}}\cdot q=0$. If $N=1+\sum_{\sigma}(N_{\sigma}n_{\sigma}-1)$, we have
@@
    Consider the setting of §4.2. Defining $\mathbb{X}_{\geqslant N}$ as above, and then $(A_{V})_{\geqslant N}$ as the sum of the graded components of $A_{V}$ whose label belongs to $\mathbb{X}_{\geqslant N}$, one checks using Lemma 4.11 and (5.9) that a graded $A_{V}$-module $Q$ is negligible if and only if for any $q\in Q$ there exists $N>0$ such that $(A_{V})_{\geqslant N}\cdot q=0$. Similar comments apply in the setting of §4.3.
@@
${\mathsf{Mod}}^{M}(A)\xrightarrow{Q\mapsto\widetilde{Q}}\mathsf{QCoh}({\operatorname{Proj}}^{M}(A))\longrightarrow{\mathsf{Mod}}({\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)})$
@@
To prove the proposition we need to define functorial morphisms $\varepsilon_{{\mathcal{Q}}}:\widetilde{\Gamma_{\bullet}({\mathcal{Q}})}\to{\mathcal{Q}}$ for ${\mathcal{Q}}\in{\mathsf{Mod}}({\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)})$ and $\eta_{Q}:Q\to\Gamma_{\bullet}(\widetilde{Q})$ for $Q\in{\mathsf{Mod}}^{M}(A)$, which satisfy the usual zigzag relations (see e.g. [Ac, Proposition A.1.16]). Here $\varepsilon$ is provided by Lemma 5.9. (This does not require any assumption.) To define $\eta$, we observe that for $Q\in{\mathsf{Mod}}^{M}(A)$ we have